OpenStudy (anonymous):

Say you have a right triangle with side a = 21 and side b = 220. What is the length of side c?

OpenStudy (ryan123345):

Use the pythagorean theorem. \[a^2 + b^2 = c^2\]

OpenStudy (ryan123345):

\[21^2 + 220^2 = c^2\]

OpenStudy (ryan123345):

\[441 + 48400 = c^2\]

OpenStudy (ryan123345):

\[48841 = c^2\]

OpenStudy (ryan123345):

\[c = \sqrt{48841}\]

OpenStudy (ryan123345):

\[c = 221\]

OpenStudy (anonymous):

That makes so much more sense now thank you (: I loved how you broke it down too it really helped, thanks again (: